7 years ago i was 11 times as old as my daughter. Now i am 4 times as old as her. How old am i?
Answer by ankor@dixie-net.com(12692)


:
Let A = my age now
Let D = daughter's age now
:
In 7 years: my age = (A-7) and daughter's age (D-7)
:
"7 years ago i was 11 times as old as my daughter".
(A-7) = 11(D-7)
A - 7 = 11D - 77
A = 11D - 77 + 7; add 7 to both sides
A = (11D - 70)
:
"Now i am 4 times as old as her."
A = 4D
:
How old am i?
Substitute (11D-70) for A in the above equation:
11D - 70 = 4D
11D - 4D = +70
7D = 70
D = 70/7
D = 10 yrs is the daughter's present age
:
Remember A = 4D
A = 4(10)
A = 40 yrs is my age now
:
:
Confirm that the statement "7 years ago i was 11 times as old as my daughter."
40 - 7 = 11(10-7)
33 = 33
:
Could you follow the reasoning on this OK?